OpenStudy (anonymous):

For the function f(x) = |2x - 7| - 3, find f(2).

OpenStudy (anonymous):

Plug in 2 for x and if and when you get a negative number inside | | it will become positive. For example: f(x) = |6x -15| + 2, find (2) = |6*2 - 15| + 2 = |12 - 15| +2 = |-3| + 2 = 3 + 2 = 5

OpenStudy (anonymous):

\[f(2)=\left| 2*2-7 \right|-3=\left| -3 \right|-3=3-3=0\]

OpenStudy (anonymous):

Yes.